OpenStudy (anonymous):

explain how to graph x + 3y = 9 using the x- and y-intercepts method.

jimthompson5910 (jim_thompson5910):

do you know how to find the x intercept?

OpenStudy (anonymous):

by looking on the right of the graph?

jimthompson5910 (jim_thompson5910):

yes, but let's say you didn't know what the graph looked like or didn't have a graphing calculator

jimthompson5910 (jim_thompson5910):

and you only had graphing paper

OpenStudy (anonymous):

ok,umm .. :( i dont know, i am not sure.

jimthompson5910 (jim_thompson5910):

Let's find the x-intercept

jimthompson5910 (jim_thompson5910):

Plug in y = 0 and solve for x x + 3y = 9 x + 3(0) = 9 x + 0 = 9 x = 9 So the x-intercept is the point (9,0)

jimthompson5910 (jim_thompson5910):

Let me know when you're ready to move on to the next part

OpenStudy (anonymous):

ok i am starting to understand that part ..:D

jimthompson5910 (jim_thompson5910):

ready for the y-intercept?

OpenStudy (anonymous):

yes :D

jimthompson5910 (jim_thompson5910):

Plug in x = 0 and solve for y x + 3y = 9 0 + 3y = 9 3y = 9 y = 9/3 y = 3 So the y-intercept is the point (0,3)

jimthompson5910 (jim_thompson5910):

Notice how everything flipped

jimthompson5910 (jim_thompson5910):

So we have the x-intercept (9,0) and the y-intercept (0,3)

OpenStudy (anonymous):

yea i noticed :D

jimthompson5910 (jim_thompson5910):

alright great

jimthompson5910 (jim_thompson5910):

with these two points, you plot them on a grid or graph paper

jimthompson5910 (jim_thompson5910):

then you draw a straight line through them to graph x+3y = 9

OpenStudy (anonymous):

omg i think i am starting to understand now :D

jimthompson5910 (jim_thompson5910):

yay, that's good lol

OpenStudy (anonymous):

lol! thank you!

jimthompson5910 (jim_thompson5910):

you're welcome